i can't solve this two problems and i been at it for 2 hours can i get some help?
first problem
n+3/3-n/4=n-2/5
second problem
x+4/3-x/7= x+7/5

Answer by ikdeep(226)

The problem which you have give is a bit confusing, I need some clarification,
In you first problem i.e n+3/3-n/4=n-2/5,,,here (3-n)is divided by 4 or n is divided my 4 (similar clarification needed where ever you put "/" sigh ),,,when two or more variables are multiplied or divided by another variable ,,please always put those variables in brackets.
is you problem looks like this (n+3)/(3-n)/4=(n-2)/5
Contact again with proper format